OpenStudy (anonymous):

c(x) = 1500 +10x +0.2x^2 (0<=x<=150) Items sell for $150 each. What is the revenue function?

jimthompson5910 (jim_thompson5910):

Revenue = (Price Per Item)*(Number of Items Sold) Revenue = ($150)*(x) Revenue = 150x So the revenue function is R(x) = 150x where x is the number of items sold (each at $150)

jimthompson5910 (jim_thompson5910):

c(x) = 1500 +10x +0.2x^2 looks like the cost function so profit would be Profit = (Revenue) - (Cost) and you can use those two functions to calculate the profit function

OpenStudy (anonymous):

I was totally over thinking this. Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome